The value A is 7

Insert YOUR value of A in the following questions (all). Question 1 (30%) For the following probability density function (p(INSERT THE VALUE OF A.) From the following Distribution Function: What is the probability to get: (2a) x < 2A (i.e., x les
Answer #1

& A=7 f( 6 aco O<31 E(X) = 52x+d2 E(+²) = /2x3 da - 2 [x4] ~CX) = E(x²) - [€(x)]? 18Fach Fra) xco 37 27,7 then pdf ir 6/7 PCX <14) - P(x=0)+P(=7) 11 PL x 276) = P(xc35) - P(x-o) 22- -